# Surfaces and other Peano Continua with no Generic Chains

### v1

@@ 5. Walks on Brick Partitions and their Refinements
The collection $\{W\subseteq X:W\in\mathcal{W}\text{ for some brick partition }\mathcal{W}\preceq\mathcal{U}\}$ is $\pi$-basis of $X$, so, up to restricting $\mathbf{V}$, we can suppose that there is a brick partition $\mathcal{V}\preceq\mathcal{U}$ and finitely many subsets $\mathcal{V}_{i}\subseteq\mathcal{V}$ so that $\mathbf{V}$ is of the form
$\mathbf{V}=\{c\in\Phi(X):c\cap O(\mathcal{V}_{i})\neq\emptyset\text{ for every}i<n\}.$
@@
Let $v$ be any reduced walk which starts in $\mathcal{V}_{0}$, and successively covers each $\mathcal{V}_{i}$, taking care of not exiting it before having visited each of its elements. Then $\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}\subseteq\mathbf{O}_{\mathcal{U},u}$, and thus $\rho^{\mathcal{V}}_{\mathcal{U}}\cdot v\sqsubseteq u$, by (5.1). ∎
We state two preliminary lemmas which will be needed later on.
Lemma 5.3.
Let $X$ be a Peano continuum, fix a brick partition $\mathcal{W}$ and a reduced walk $w$ on $\mathcal{W}$. For any brick partition $\mathcal{V}\preceq\mathcal{W}$, there is a reduced walk $v$ on $\mathcal{V}$ such that $\rho^{\mathcal{V}}_{\mathcal{U}}\cdot v$ monotonically refines $w$. Moreover, $v$ can be chosen to be a path and/or circular/lasso, whenever $w$ is.
Proof.
Let $P_{0}$ be a path of length $1$ in $\mathcal{V}(w(0))$, and choose it moreover to be $E$-adjacent to $\mathcal{V}(w(\operatorname{lh}(w)-1))$ if $w$ is circular. Suppose we have defined $P_{i}$, for $i<\operatorname{lh}(w)-1$, such that $P_{i}(-1)\in\mathcal{V}(w(i))$, and find a path $P_{i+1}$ from $P_{i}(-1)$ to $\mathcal{V}(w(i+1))$ in $\mathcal{V}(w(i))\cup\mathcal{V}(w(i+1))$. Notice that all of $P_{i+1}$ except its last element is in $\mathcal{V}(w(i))$.
Adapt the last step accordingly if $w$ is a lasso or circular: in this latter case set the only element of $P_{0}$ to be the final endpoint of $P_{i+1}$. The reduced concatenation of the $P_{i}$’s is the required walk $v$. If $w$ is a path, then the $P_{i}$’s meet at most in their endpoints, so $v$ is again a path. ∎
@@
[…]
@@ 5.1. A Criterion for the Absence of Generic Chains
We recall the following criterion due to Rosendal for verifying the existence of comeager orbits. Recall that a continuous action $H\curvearrowright Y$ is topologically transitive if $H\mathbf{U}\cap\mathbf{V}\neq\emptyset$ for any nonempty open sets $\mathbf{U},\mathbf{V}\subseteq Y$.
Theorem 5.5 ([ben2017metrizable]*Proposition 3.2).
@@
The goal of this section is to obtain a combinatorial formulation of (1) $\Rightarrow$ (2) from Theorem 5.5 when $Y=\Phi(X)$ is the space of maximal connected chains over a Peano continuum $X$ and $H=\operatorname{Homeo}(X)$ is its homeomorphism group. This is done in Theorem 5.8, after establishing our discretization theorem for homeomorphism of a Peano continuum, Theorem 5.7. The following is a generalization of [MOT]*Corollary 2.11 (where $\mathcal{U}=\mathcal{V}$).
@@
Up to taking a refinement we can assume that $\mathcal{U}$ is as fine as desired. For each $V\in\mathcal{V}$, let $T_{V}$ be a continuum theoretical tree in $V$ which meets every $U\in\operatorname{Star}(V,\mathcal{U})$. Let $\mathcal{W}_{0}\preceq\mathcal{U}$ be a brick partition fine enough so that $\operatorname{Star}(T_{V},\mathcal{W}_{0})\cap\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W}_{0})=\emptyset$ for all $V\in\mathcal{V}$.
@@
Indeed, suppose that there is $W\in\mathcal{B}^{U}_{i}\cap\operatorname{Core}(V,\mathcal{W}_{0})$. Since $\mathcal{W}_{0}(U)$ is connected and $U\cap V\neq\emptyset$, there is a path $P$ from $W\in\mathcal{B}^{U}_{i}$ to $\operatorname{Star}(T_{V},\mathcal{W}_{0})$ in $\mathcal{W}_{0}(U)$. By assumption $W\not\in\mathcal{C}^{U,V}$, so there must be a minimal $k<\operatorname{lh}(P)$ such that $P(k)\in\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W}_{0})$. For each $j<k$, $P(j)\in\mathcal{W}_{0}(U\cap V)$, so $P(j)\not\in\mathcal{C}^{U,V}$, as otherwise there would be a path from $W$ to $\operatorname{Star}(T_{V},\mathcal{W}_{0})$ in $\mathcal{W}_{0}(U\cap V)$ which avoids $\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W}_{0})$. But $P(j)\not\in\mathcal{C}^{U,V^{\prime}}$, for $V^{\prime}\neq V$, because it is contained in $V$, so $P(j)\in\mathcal{B}^{U}$. Moreover, $P(k)\in\mathcal{B}^{U}$, since it belongs to $\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W}_{0})$. Since $\mathcal{B}^{U}_{i}$ is a connected component, $P(k)\in\mathcal{B}^{U}_{i}$, so $\mathcal{B}^{U}_{i}\nsubseteq\operatorname{Core}(V,\mathcal{W}_{0})$.
It follows that
@@
Pick $\mathcal{W}$ which is moreover fine enough that for each $s<\operatorname{lh}(u)$ and $t<\operatorname{lh}(v)$:
@@
for all $s<\operatorname{lh}(u)$, so $\rho^{\mathcal{W}}_{\mathcal{U}}\cdot w\sqsubseteq u$. Fix $t<\operatorname{lh}(v)$. By (a) and (5.5),
$\alpha[\{w(0),\dots,w(k_{t})\}]=\alpha[\operatorname{Star}(C_{t}^{v},\mathcal{W})]\subseteq\{(v(0)),\dots,(v(t))\}.$
@@
$\alpha[\{w(0),\dots,w(k_{t})\}]=\{(v(0)),\dots,(v(t))\}.$
@@
$d_{\mathcal{U}}(\rho^{\mathcal{U}_{0}}_{\mathcal{U}}(U),\rho^{\mathcal{U}_{0}}_{\mathcal{U}}(U^{\prime}))\leq 1\text{ whenever }d_{\mathcal{U}_{0}}(U,U^{\prime})\leq 2.$
Find a symmetric open neighborhood of the identity $H_{0}\subseteq H_{\mathcal{U}_{0},1}$ (see (5.6)). Also, set $\mathbf{U}\coloneqq\mathbf{O}_{\mathcal{U}_{0},u_{0}}$. Theorem 5.5 gives us a nonempty $\mathbf{V}\subseteq\mathbf{U}$. By Proposition 5.2, there are $(\mathcal{V},v)\preceq(\mathcal{U}_{0},u_{0})\preceq(\mathcal{U},u)$, with $\mathcal{V}$ a brick partition, such that $\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}$.
Suppose now that we are given a brick partition $\mathcal{W}\preceq\mathcal{V}$ and two walks $w_{0},w_{1}$ on $\mathcal{W}$ with $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w_{i}\sqsubseteq v$, for $i=0,1$. Set $\mathbf{W}_{i}\coloneqq\mathbf{O}_{\mathcal{W},w_{i}}$; then $\mathbf{W}_{i}\subseteq\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}$, for $i=0,1$. Therefore by Theorem 5.5 there is $h\in H_{0}$ such that $h^{-1}\cdot\mathbf{W}_{0}\cap\mathbf{W}_{1}$ is a nonempty open set. We can thus apply Proposition 5.2 to obtain $(\mathcal{Z}^{\prime},z^{\prime})\preceq(\mathcal{W},w_{1})$ such that $\mathbf{O}_{\mathcal{Z}^{\prime},z^{\prime}}\subseteq h^{-1}\cdot\mathbf{W}_{0}\cap\mathbf{W}_{1}\subseteq\mathbf{W}_{1}$.
@@ 6. Absence of Generic Chains
In this section we finally put all pieces together and prove Theorem 1.2. The first step is to show that the hypothesis that $X$ has no locally separating points allows us to only deal with walks which are spaced paths. We do so in Theorem 6.2 ( (Upgrading Walks to Paths).). This will be helpful when trying to use Lemma 3.2 and Lemma 3.3 to compare winding numbers of walks defined on different partitions.
@@
Let $X$ be a Peano continuum with no locally separating points. Let $\mathcal{U}$ be a brick partition of $X$ and $u$ be a path on $\mathcal{U}$. There are a brick partition $\mathcal{V}\preceq\mathcal{U}$ and $2$ paths $v_{0},v_{1}$ on $\mathcal{V}$ which are disjoint except for their endpoints and such that $\rho^{\mathcal{V}}_{\mathcal{U}}\cdot v_{i}$ monotonically refines $u$ for $i=0,1$.
@@
For each $i<\operatorname{lh}(u)-1$, let $A_{i}$ be a subset of cardinality 2 of $\operatorname{Cl}(u(i))\cap\operatorname{Cl}(u(i+1))\cap(u(i)\vee u(i+1))$, which exists by 4.2.(2). Given $11<i<\operatorname{lh}(u)-$ 1, 4.3.(2) gives two disjoint arcs $\gamma_{i},\gamma^{\prime}_{i}$ in $u(i)\cup A_{i}\cup A_{i+1}$ connecting $A_{i}$ to $A_{i+1}$. Since $u$ is a path, all the $A_{i}$’s and all the $\gamma_{i}$’s are distinct, and the $\gamma_{i}$’s intersect at most in their endpoints. The same holds for the $\gamma^{\prime}_{i}$’s. Let $\lambda,\lambda^{\prime}$ be the two disjoint arcs resulting from the concatenations of the $\gamma_{i}$’s and $\gamma^{\prime}_{i}$’s respectively.
Let $\tilde{\mathcal{V}}\preceq\mathcal{U}$ be a brick partition fine enough that $\operatorname{Star}(\lambda,\tilde{\mathcal{V}})\cap\operatorname{Star}(\lambda^{\prime},\tilde{\mathcal{V}})=\emptyset$. Let, for each $11<i<\operatorname{lh}(u)-$ 1:
@@
Let $\mathcal{V}$ be an amalgam of $\tilde{\mathcal{V}}$ refining $\mathcal{U}$ and containing each $\bigvee P_{i}$ and each $\bigvee P^{\prime}_{i}$, for $i<\operatorname{lh}(u)$. Then $v_{0}\coloneqq\langle P_{0},\dots,P_{\operatorname{lh}(u)-1}\rangle$ and $v_{1}\coloneqq\langle P^{\prime}_{0},\dots,P^{\prime}_{\operatorname{lh}(u)-1}\rangle$ are paths on $\mathcal{V}$ satisfying the conclusions of the proposition. ∎
@@
Let $X$ be a Peano continuum with no locally separating points. Let $\mathcal{V}$ be a brick partition of $X$ and $v$ be a reduced walk on $\mathcal{V}$. There are a spaced path $w$ on a brick partition $\mathcal{W}$ such that $(\mathcal{W},w)\preceq(\mathcal{V},v)$. If furthermore $v$ is uncrossed then $w$ can be chosen so that $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w$ monotonically refines $v$.
@@
We proceed by induction on the length $\ell$ of $v$. If $\ell=1$ let $\mathcal{W}\coloneqq\mathcal{V}$ and $w\coloneqq v$. Suppose the result is true for $\ell\geq 1$ and let us prove it for $v=\langle v(0),\dots,v(\ell)\rangle$. By the induction hypothesis, there is a brick partition $\mathcal{Z}\preceq\mathcal{V}$ and a spaced path $z$ on $\mathcal{Z}$ such that $\rho^{\mathcal{Z}}_{\mathcal{V}}\cdot z$ refines $\langle v(0),\dots,v(\ell-1)\rangle$, and it does so monotonically if the latter is uncrossed.

### v3

@@ 5. Walks on Brick Partitions and their Refinements
Set $\varepsilon\coloneqq\mathrm{mesh}_{d}(\mathcal{U})$. Recall that in order to show that $d_{\Phi}(\mathcal{C},\mathcal{C}^{\prime})\leq\varepsilon$, it suffices to prove that for every $K\in\mathcal{C}$ there is $K^{\prime}\in\mathcal{C}^{\prime}$ such that $d_{H}(K,K^{\prime})\leq\varepsilon$, and symmetrically that for every $K^{\prime}\in\mathcal{C}^{\prime}$ there is $K\in\mathcal{C}$ such that $d_{H}(K,K^{\prime})\leq\varepsilon$.
To this end, fix $K\in\mathcal{C}$. By hypothesis, there is $K^{\prime}\in\mathcal{C}^{\prime}$ such that $\operatorname{Star}(K,\mathcal{U})=\operatorname{Star}(K^{\prime},\mathcal{U})$. It follows that $d_{H}(K,K^{\prime})\leq\varepsilon$.
Fix next $K^{\prime}\in\mathcal{C}^{\prime}$ and let $K\in\mathcal{C}$ be the minimum element such that $\operatorname{Star}(K^{\prime},\mathcal{U})\subseteq\operatorname{Star}(K,\mathcal{U})$. If $K$ is the root of $\mathcal{C}$, then
$K\subseteq\bigcap\{\operatorname{Cl}(U):U\in\operatorname{Star}(K,\mathcal{U})\}\subseteq\bigcap\{\operatorname{Cl}(U):U\in\operatorname{Star}(K^{\prime},\mathcal{U})\},$
which implies $d_{H}(K,K^{\prime})\leq\varepsilon$. Otherwise, by maximality and connectedness of $\mathcal{C}$ (as a subset of $C(X)$, see [illanes1999hyperspaces]*Lemma 14.7), see , we have that
$K=\operatorname{Cl}\left(\bigcup\{F\in\mathcal{C}:F\subsetneq K\}\right).$
The minimality assumption on $K$ and (5.3) entail $\operatorname{Star}(F,\mathcal{U})\subsetneq\operatorname{Star}(K^{\prime},\mathcal{U})$ for every $F\in\mathcal{C}$ with $F\subsetneq K$, hence we have that $K\subseteq\operatorname{Cl}(\bigcup\operatorname{Star}(K^{\prime},\mathcal{U}))$. This, along with $\operatorname{Star}(K^{\prime},\mathcal{U})\subseteq\operatorname{Star}(K,\mathcal{U})$, is sufficient to deduce $d_{H}(K,K^{\prime})\leq\varepsilon$. ∎
Proof of Proposition 5.2.
Let $\varepsilon\coloneqq\mathrm{mesh}_{d}(\mathcal{W})$. Given a walk $w$ on $\mathcal{W}$ such that $\operatorname{Im}(w)=\mathcal{W}$, any two chains $\mathcal{C},\mathcal{C}^{\prime}\in\mathbf{O}_{\mathcal{W},w}$ satisfy
$\{\operatorname{Star}(K,\mathcal{U}):K\in\mathcal{C}\}=\{\operatorname{Star}(K^{\prime},\mathcal{U}):K^{\prime}\in\mathcal{C}^{\prime}\}.$
Lemma 5.3 then shows that the $d_{\Phi}$-diameter of each $\mathbf{O}_{\mathcal{W},w}\in\Phi(\mathcal{W})$ is at most $\varepsilon$.
The moreover statement follows from (5.1). We can thus use it to prove that $\bigcup\Phi(\mathcal{W})$ is dense in $X$: more precisely, since $\mathcal{W}$ can be refined by arbitrarily fine brick partitions, it suffices to show that $\bigcup\Phi(\mathcal{W})$ is $\varepsilon$-dense. Fix some $\mathcal{C}\in\Phi(X)$. The set $\{\operatorname{Star}(K,\mathcal{W}):K\in\mathcal{C}\}$ consists of a finite number of $E$-connected subsets of $\mathcal{W}$ linearly ordered by inclusion
$\mathcal{W}_{0}\subset\cdots\subset\mathcal{W}_{m-1}=\mathcal{W}.$
Let $w$ be any walk which starts in $\mathcal{W}_{0}$, and successively covers each $\mathcal{W}_{i}$, taking care of not exiting it before having visited each of its elements. By Lemma 5.3, for any $\mathcal{C}^{\prime}\in\mathbf{O}_{\mathcal{W},w}$, one has $d_{\Phi}(\mathcal{C},\mathcal{C}^{\prime})\leq\varepsilon$.
Corollary 5.4.
Let $X$ be a Peano continuum, and fix a brick partition $\mathcal{U}$ and a walk $u$ on $\mathcal{U}$. Let $\mathbf{V}\subseteq\mathbf{O}_{\mathcal{U},u}\subseteq\Phi(X)$ be a nonempty open set. There exist a brick partition and a reduced walk $(\mathcal{V},v)\preceq(\mathcal{U},u)$ such that $\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}$.
Proof.
Up to passing to a smaller open set, we can suppose that $\mathbf{V}$ is the $d_{\Phi}$-ball of radius $\varepsilon$ centered around some $\mathcal{C}\in\mathbf{O}_{\mathcal{U},u}$. Let $\mathcal{V}\preceq\mathcal{U}$ be a brick partition whose mesh is smaller than $\varepsilon$. By Proposition 5.2, there is a reduced walk $v$ on $\mathcal{V}$ such that $\operatorname{Im}(v)=\mathcal{V}$ and $\mathcal{C}\in\operatorname{Cl}(\mathbf{O}_{\mathcal{V},v})$. By the same proposition, the diameter of $\mathbf{O}_{\mathcal{V},v}$ is smaller than $\varepsilon$, so $\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}\subseteq\mathbf{O}_{\mathcal{U},u}$. Since $\operatorname{Im}(v)=\mathcal{V}$, we can apply (5.2) and conclude that $\rho^{\mathcal{V}}_{\mathcal{U}}\cdot v\sqsubseteq u$. ∎
@@
Let $X$ be a Peano continuum, and fix a brick partition $\mathcal{W}$ and a path $w$ on $\mathcal{W}$. There are a brick partition $\mathcal{Z}\preceq\mathcal{W}$ and a spaced path $z$ on $\mathcal{Z}$ such that $\rho^{\mathcal{Z}}_{\mathcal{W}}\cdot z$ monotonically refines $w$ and $\rho^{\mathcal{Z}}_{\mathcal{W}}([z]_{1})\subseteq\operatorname{Im}(w)$. If $w$ is moreover circular, $z$ can also be chosen to be circular.
@@ 5.1. A Criterion for the Absence of Generic Chains
We recall the following criterion due to Rosendal for verifying the existence of comeager orbits, whose proof appears in [ben2017metrizable]*Proposition 3.2. Recall that a continuous action $H\curvearrowright Y$ is topologically transitive if $H\cdot\mathbf{U}\cap\mathbf{V}\neq\emptyset$ for any nonempty open sets $\mathbf{U},\mathbf{V}\subseteq Y$.
Theorem 5.6 (Rosendal’s criterion).
@@
The goal of this section is to obtain a combinatorial formulation of (1) $\Rightarrow$ (2) from (Rosendal’s criterion). (Theorem 5.6) when $Y=\Phi(X)$ is the space of maximal connected chains on a Peano continuum $X$ and $H=\operatorname{Homeo}(X)$ is its homeomorphism group. This is done in Theorem 5.9, after establishing our discretization theorem for homeomorphism of a Peano continuum, Theorem 5.8. The following is a generalization of [MOT]*Corollary 2.11 (where $\mathcal{U}=\mathcal{V}$).
@@
Up to taking a refinement we can assume that $\mathcal{U}$ is already as fine as desired. For each $V\in\mathcal{V}$, let $T_{V}$ be a continuum-theoretical tree in $V$ which meets every $U\in\operatorname{Star}(V,\mathcal{U})$. Let $\mathcal{W}_{0}\preceq\mathcal{U}$ be a brick partition fine enough so that $\operatorname{Star}(T_{V},\mathcal{W}_{0})\cap\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W}_{0})=\emptyset$ for all $V\in\mathcal{V}$.
@@
Indeed, suppose that there is $W\in\mathcal{B}^{U}_{i}\cap\operatorname{Core}(V,\mathcal{W}_{0})$. Since $\mathcal{W}_{0}(U)$ is $E$-connected and $U\cap V\neq\emptyset$, there is a path $P$ from $W\in\mathcal{B}^{U}_{i}$ to $\operatorname{Star}(T_{V},\mathcal{W}_{0})$ in $\mathcal{W}_{0}(U)$. By assumption $W\not\in\mathcal{C}^{U,V}$, so there must be a minimal $k<\operatorname{lh}(P)$ such that $P(k)\in\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W}_{0})$. For each $j<k$, $P(j)\in\mathcal{W}_{0}(U\cap V)$, so $P(j)\not\in\mathcal{C}^{U,V}$, as otherwise there would be a path from $W$ to $\operatorname{Star}(T_{V},\mathcal{W}_{0})$ in $\mathcal{W}_{0}(U\cap V)$ which avoids $\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W}_{0})$. But $P(j)\not\in\mathcal{C}^{U,V^{\prime}}$, for $V^{\prime}\neq V$, because it is contained in $V$, so $P(j)\in\mathcal{B}^{U}$. Moreover, $P(k)\in\mathcal{B}^{U}$, since it belongs to $\operatorname{Star}(\operatorname{Bd}(V),\mathcal{W}_{0})$. Since $\mathcal{B}^{U}_{i}$ is a connected component, $P(k)\in\mathcal{B}^{U}_{i}$, so $\mathcal{B}^{U}_{i}\nsubseteq\operatorname{Core}(V,\mathcal{W}_{0})$.
Notice that
@@
Pick $\mathcal{W}$ which is moreover fine enough that for each $t<\operatorname{lh}(v)$:
@@
for all $s<\operatorname{lh}(u)$, so $\rho^{\mathcal{W}}_{\mathcal{U}}\cdot w\sqsubseteq u$, since $\operatorname{Im}(\rho^{\mathcal{W}}_{\mathcal{U}}\cdot w)=\rho^{\mathcal{W}}_{\mathcal{U}}[\operatorname{Im}(w)]\supseteq\operatorname{Im}(u)$. Fix $t<\operatorname{lh}(v)$. By (a) and (5.7),
$\alpha[\{w(0),\dots,w(k_{t})\}]=\alpha[\operatorname{Star}(C_{t}^{v},\mathcal{W})]\subseteq\{v(0),\dots,v(t)\}.$
@@
$\alpha[\{w(0),\dots,w(k_{t})\}]=\{v(0),\dots,v(t)\}.$
@@
(5.9) $d_{\mathcal{U}}(\rho^{\mathcal{U}_{0}}_{\mathcal{U}}(U),\rho^{\mathcal{U}_{0}}_{\mathcal{U}}(U^{\prime}))\leq 1\text{ whenever }d_{\mathcal{U}_{0}}(U,U^{\prime})\leq 2.$
Find a symmetric open neighborhood of the identity $H_{0}\subseteq H_{\mathcal{U}_{0},1}$ (see (5.8)). Also, set $\mathbf{U}\coloneqq\mathbf{O}_{\mathcal{U}_{0},u_{0}}$. (Rosendal’s criterion). (Theorem 5.6) gives us a nonempty $\mathbf{V}\subseteq\mathbf{U}$. By Corollary 5.4, there are $(\mathcal{V},v)\preceq(\mathcal{U}_{0},u_{0})\preceq(\mathcal{U},u)$, with $\mathcal{V}$ a brick partition, such that $\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}$.
Suppose now that we are given a brick partition $\mathcal{W}\preceq\mathcal{V}$ and two walks $w_{0},w_{1}$ on $\mathcal{W}$ with $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w_{i}\sqsubseteq v$, for $i=0,1$. Set $\mathbf{W}_{i}\coloneqq\mathbf{O}_{\mathcal{W},w_{i}}$; then $\mathbf{W}_{i}\subseteq\mathbf{O}_{\mathcal{V},v}\subseteq\mathbf{V}$, for $i=0,1$. Therefore by (Rosendal’s criterion). there is $h\in H_{0}$ such that $h^{-1}\cdot\mathbf{W}_{0}\cap\mathbf{W}_{1}$ is a nonempty open set. We can thus apply Corollary 5.4 to obtain $(\mathcal{Z}^{\prime},z^{\prime})\preceq(\mathcal{W},w_{1})$ such that $\mathbf{O}_{\mathcal{Z}^{\prime},z^{\prime}}\subseteq h^{-1}\cdot\mathbf{W}_{0}\cap\mathbf{W}_{1}\subseteq\mathbf{W}_{1}$.
@@ 6. Absence of Generic Chains
In this section we finally put all pieces together and prove Theorem 1.2.
6.1. Walks on Peano continua with no locally separating points
The first step towards Theorem 1.2 is to show that the hypothesis that $X$ has no locally separating points allows us to only deal with walks which are spaced paths. We do so in Theorem 6.2 ( (Upgrading Walks to Paths).). This will be helpful when trying to use Lemma 3.3 and Lemma 3.4 to compare winding numbers of walks defined on different partitions.
@@ 6.1. Walks on Peano continua with no locally separating points
Let $X$ be a Peano continuum with no locally separating points. Let $\mathcal{U}$ be a brick partition of $X$ and $u$ be a path on $\mathcal{U}$. There are a brick partition $\mathcal{V}\preceq\mathcal{U}$ and two paths $v_{0},v_{1}$ on $\mathcal{V}$ which have the same endpoints and are disjoint elsewhere and such that $\rho^{\mathcal{V}}_{\mathcal{U}}\cdot v_{i}$ monotonically refines $u$ for $i=0,1$.
@@
For each $i<\operatorname{lh}(u)-1$, let $A_{i}$ be a subset of cardinality 2 of $\operatorname{Cl}(u(i))\cap\operatorname{Cl}(u(i+1))\cap(u(i)\vee u(i+1))$, which exists by 4.2.(2). Given $11\leq i<\operatorname{lh}(u)-$ 1, 4.3.(2) gives two disjoint arcs $\gamma_{i},\gamma^{\prime}_{i}$ in $u(i)\cup A_{i-1}\cup A_{i}$ connecting $A_{i-1}$ to $A_{i}$. Since $u$ is a path, all the $A_{i}$’s and all the $\gamma_{i}$’s are distinct, and the $\gamma_{i}$’s intersect at most in their endpoints. The same holds for the $\gamma^{\prime}_{i}$’s. Let $\lambda,\lambda^{\prime}$ be the two disjoint arcs resulting from the concatenations of the $\gamma_{i}$’s and $\gamma^{\prime}_{i}$’s respectively.
Let $\tilde{\mathcal{V}}\preceq\mathcal{U}$ be a brick partition fine enough that $\operatorname{Star}(\lambda,\tilde{\mathcal{V}})\cap\operatorname{Star}(\lambda^{\prime},\tilde{\mathcal{V}})=\emptyset$. Let, for each $11\leq i<\operatorname{lh}(u)-$ 1:
@@
Let $\mathcal{V}$ be an amalgam of $\tilde{\mathcal{V}}$ refining $\mathcal{U}$ and containing each $\bigvee P_{i}$ and each $\bigvee P^{\prime}_{i}$, for $i<\operatorname{lh}(u)$. Then $v_{0}\coloneqq\langle\bigvee P_{0},\dots,\bigvee P_{\operatorname{lh}(u)-1}\rangle$ and $v_{1}\coloneqq\langle\bigvee P^{\prime}_{0},\dots,\bigvee P^{\prime}_{\operatorname{lh}(u)-1}\rangle$ are paths on $\mathcal{V}$ satisfying the conclusions of the proposition. ∎
@@
Let $X$ be a Peano continuum with no locally separating points. Let $\mathcal{V}$ be a brick partition of $X$ and $v$ be a reduced walk on $\mathcal{V}$. There is a spaced path $w$ on a brick partition $\mathcal{W}$ such that $(\mathcal{W},w)\preceq(\mathcal{V},v)$. If furthermore $v$ is uncrossed then $w$ can be chosen so that $\rho^{\mathcal{W}}_{\mathcal{V}}\cdot w$ monotonically refines $v$.
@@
We proceed by induction on the length $\ell$ of $v$. If $\ell=1$, let $\mathcal{W}\coloneqq\mathcal{V}$ and $w\coloneqq v$. Suppose the result is true for $\ell\geq 1$ and let us prove it for $v=\langle v(0),\dots,v(\ell)\rangle$. By the induction hypothesis, there is a brick partition $\mathcal{Z}\preceq\mathcal{V}$ and a spaced path $z$ on $\mathcal{Z}$ such that $\rho^{\mathcal{Z}}_{\mathcal{V}}\cdot z$ refines $\langle v(0),\dots,v(\ell-1)\rangle$, is confined to it, $\rho^{\mathcal{Z}}_{\mathcal{V}}(z(-1))=v(\ell-1)$, and the refinement is monotonic if the latter is uncrossed.